OpenStudy (anonymous):

Solve the equation. -6p - 21 = 3p - 12

OpenStudy (anonymous):

-6p -21 = 3p - 12 +21 +21 -6p = 3p +9 -3p -3p -9p=9 solve for p

OpenStudy (anonymous):

Thankkk youuuu!

OpenStudy (anonymous):

x/5 + 9 = 4 do you know this one? @larryboxaplenty

OpenStudy (anonymous):

(x/5) + 9 = 4 -9 -9 (x/5) = -5 solve for x

OpenStudy (anonymous):

These are my answers @larryboxaplenty A. 65 B. –25 C. 5 D. 20

OpenStudy (anonymous):

Would it be positive 5?

OpenStudy (anonymous):

since the two negatives make a positive?

OpenStudy (anonymous):

(x/5) = -5 times each side by 5 x = (-5*5) x = -25 B